# Distributionally Robust Markov Games with Average Reward

### v1

@@ Abstract
This paper introduces the formulation of a distributionally robust Markov game (DR-MG) with average rewards, a crucial framework for multi-agent decision-making under uncertainty over extended horizons. Unlike finite-horizon or discounted models, the average-reward criterion naturally captures long-term performance for systems designed for continuous operation, where sustained reliability is paramount. We account for uncertainty in transition kernels, with players aiming to optimize their worst-case average reward. We first establish a connection between the multi-agent and single agent settings, and derive the solvability of the robust Bellman equation under the average-reward formulation. We then rigorously prove the existence of a robust Nash Equilibrium (NE), offering essential theoretical guarantees for system stability. We further develop and analyze an algorithm named robust Nash-Iteration to compute the robust Nash Equilibria among all agents, providing practical tools for identifying optimal strategies in complex, uncertain, and long-running multi-player environments. Finally, we demonstrate the connection between the average-reward NE and the well-studied discounted NEs, showing that the former can be approximated as the discount factor approaches one. Together, these contributions provide a comprehensive theoretical and algorithmic foundation for identifying optimal strategies in complex, uncertain, and long-running multi-player environments, which allow for the future extension of robust average-reward single-agent problems to the multi-agent setting.
@@ 1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments Shapley (1953); Filar and Vrieze (2012); Littman (1994). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the "Sim-to-Real" gap McMahan, Artiglio, and Xie (2024), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks Bukharin et al. (2023); Ma et al. (2023). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
@@
To address this vulnerability, the framework of distributionally robust Markov Games (DR-MGs) has been developed Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011), extending the principles of distributionally robust Markov Decision Processes (MDPs) Bagnell, Ng, and Schneider (2001); Nilim and El Ghaoui (2004); Iyengar (2005) to the multi-agent setting. Instead of relying on a single, fixed MG model, the robust approach seeks to find an equilibrium that optimizes the worst-case performance over a predefined uncertainty set of plausible game models under potential model mismatches. The resulting robust equilibrium provides a performance guarantee across all models within this set, thereby ensuring each agent’s resilience against model mismatch.
@@
The nature of a DR-MG problem is fundamentally shaped by the optimality criterion used. Much of the prior work has concentrated on the finite-horizon Ma et al. (2023); Shi et al. (2024a, b); Jiao and Li (2024); Li et al. (2025); Blanchet et al. (2023) or discounted-reward setting Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011), where the objective is to maximize the expected reward over a finite number of steps or a discounted sum of rewards.
@@
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics Krnjaic et al. (2024), communication networks Qu et al. (2020), autonomous vehicle coordination Peake et al. (2020), financial markets Vadori et al. (2024), or even peer-to-peer energy trading Qiu et al. (2021). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate Hernández-Lerma and Lasserre (2000). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium Filar and Vrieze (2012); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions Puterman (1994); Atia et al. (2021). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
@@
Moreover, none of the methods for DR-MGs under the discounted reward or finite horizon can be extended to the average-reward setting. The existence of (non-stationary) Nash Equilibrium under finite horizon is derived through backward induction Blanchet et al. (2023), which does not apply to infinite-horizon settings. The existence of stationary Nash Equilibrium under discounted reward is derived in Kardeş, Ordóñez, and Hall (2011), which however heavily relies on the uniqueness of the solution to the discounted robust Bellman equation Iyengar (2005) and hence cannot be applied to the average reward. Another natural attempt is to use the discounted reward as an intermediate step to study the robust Nash Equilibrium under the average reward by setting the discount factor approaching to $1$. Although this limit method is effective in the single-agent setting Wang et al. (2023a), it cannot solve the DR-MGs, due to the complicated structure (also see our detailed discussion in Section 3). To date, the literature on average-reward DR-MGs remains sparse, leaving fundamental questions about the structure of robust equilibria and the design of convergent solution algorithms largely unanswered.
@@ 1.1 Contributions
In this paper, we develop a comprehensive study of DR-MGs under the average-reward formulation. Our contributions are summarized as follows. All of our studies are developed under a standard irreducibility assumption (Assumption 4.1), and can be further extended to unichain settings (see Section F in the Appendix).
Solvability of the Robust Bellman Equation: As a foundational step, we first revisit the single-agent distributionally robust average-reward MDPs. Despite existing studies all utilizing the robust Bellman equation Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025), its solvability is not proved, but rather implicitly assumed. We establish the well-posedness of the robust average-reward Bellman equation, study the structure of its solutions, and connect them with the optimal robust policies. This analysis provides structural insights and tools that are essential for solving multi-agent Markov games when long-term performance is of interest.
Existence and Algorithmic Identification of Robust Nash Equilibrium: We further introduce a formulation of induced distributionally robust MDPs to reduce the Nash Equilibrium identification in comprehensive multi-agent games to optimizing single-agent distributionally robust MDPs. We then utilize the connection revealed to derive the convexity and the semi-continuity of the optimal policy set, based on which we further establish the existence of a stationary Nash Equilibrium. This result addresses a fundamental open question, demonstrating that stable solutions exist in these complex, robust, long-horizon competitive environments.
@@ 1.2 Related Works
Non-robust Markov Games. Markov Games Littman (1994) provide the foundational mathematical framework for multi-agent sequential decision-making. The majority of early work focused on the discounted-reward setting, where the existence of a stationary Nash Equilibrium was established in Shapley (1953). The average-reward criterion, while more suitable for systems with long operational horizons, presents greater analytical challenges. Unlike the discounted case, a stationary NE is not guaranteed to exist in a general average-reward MG. Its existence has only been proven under additional structural assumptions on the dynamics of the game Guo and Yang (2008); Hernández-Lerma and Lasserre (2000); Guo and Hernández-Lerma (2003); Zheng and Guo (2024); Nowak (1999); Iwase, Tanaka, and Wakuta (1976); Küenle and Schurath (2003); Tanaka and Wakuta (1977); Jaśkiewicz and Nowak (2001); Filar and Vrieze (2012).
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, the distributionally robust MDPs are first developed in Iyengar (2005); Nilim and El Ghaoui (2004) for both finite and discounted settings. The studies under the average reward setting are recently developed in Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025); Grand-Clement, Petrik, and Vieille (2023); Grand-Clément and Petrik (2023); Chatterjee et al. (2023). While the robust Bellman equation is central to these methods, its solvability was often implicitly assumed without rigorous proof.
Distributionally Robust Markov Games. Extending distributional robustness to the multi-agent setting is a recent and active area of research. However, the literature has overwhelmingly concentrated on the finite-horizon Ma et al. (2023); Shi et al. (2024a, b); Jiao and Li (2024); Li et al. (2025); Blanchet et al. (2023) and discounted-reward criteria Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011). These settings are often more analytically tractable because the influence of future rewards diminishes, simplifying the analysis. However, none of these methods can be extended to the average-reward setting, as we shall discuss later.
@@ 2 Preliminaries and Problem Formulation
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (Littman, 1994), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\otimes_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}^{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the actions of all agents. At some state $s$, all agents act simultaneously by taking joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
@@
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on the Nash Equilibrium (NE). More specifically, a NE is a product policy $\pi^{*}=(\pi^{*}_{1},...,\pi^{*}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{*}$:
$V^{\pi^{*}}_{\mathsf{P},i}(s)\geq V^{(\pi^{*}_{-i},\pi_{i})}_{\mathsf{P},i}(s),\forall i\in\mathcal{N},\forall\pi_{i}\in\Pi_{i},s\in\mathcal{S}$ (1)
@@
where $\Pi_{i}$ denotes the set of stationary policies for agent $i$, $\pi^{*}_{-i}$ denotes the joint policy of all other agents except $i$ from $\pi^{*}$, and $(\pi^{*}_{-i},\pi_{i})$ is the joint policy of $\pi^{*}_{-i}$ and an arbitrary policy $\pi_{i}$ of agent $i$. The existence of stationary NE is derived for MGs with discounted reward Shapley (1953) and average-reward MGs with additional structural assumptions Filar and Vrieze (2012).
@@
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\otimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ Iyengar (2005); Nilim and El Ghaoui (2004); Zhang et al. (2020); Blanchet et al. (2023); Shi et al. (2024b), where the uncertainty set is independently defined over all $(s,a)$ pairs.
@@
After all agents take their actions $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}=\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}$, and robust average reward $g^{\pi}_{\mathcal{P},i}=\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{*}$ such that
$g^{\pi^{*}}_{\mathcal{P},i}(s)\geq g^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i}(s),\forall i\in\mathcal{N},\forall\pi_{i},\forall s\in\mathcal{S}.$ (2)

### v2

@@ Abstract
We study distributionally robust Markov games (DR‑MGs) with the average‑reward criterion, a crucial framework for multi-agent decision-making under uncertainty over extended horizons. We first establish a connection between the best-response policies and the optimal policies for the induced single‑agent problems. Under a standard irreducible assumption, we derive a correspondence between the optimal policies and the solutions of the robust Bellman equation, and derive the existence of stationary Nash Equilibrium (NE) based on these results. We also study a more general weakly communicating setting. We construct a set-valued map and show its value is a subset of the best-response policies, convex and upper hemi-continuous, which imply the existence of NE. We then introduce Robust Nash‑Iteration, and provide convergence guarantees. Finally, we connect average‑reward NE to discounted robust equilibria, showing approximation as the discount factor approaches one. Our studies provide comprehensive theoretical and algorithmic foundation for decision-making in complex, uncertain, and long-running multi-player environments.
@@ 1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments (shapley1953stochastic; filar2012competitive; littman1994markov). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the "Sim-to-Real" gap (mcmahan2024roping), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks (bukharin2023robust; ma2023decentralized). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
@@
To address this vulnerability, the framework of distributionally robust Markov Games (DR-MGs) has been developed (zhang2020robust; kardecs2011discounted), extending the principles of distributionally robust Markov Decision Processes (MDPs) (bagnell2001solving; nilim2004robustness; iyengar2005robust) to the multi-agent setting. Instead of relying on a single, fixed MG model, the robust approach seeks to find an equilibrium that optimizes the worst-case performance over a predefined uncertainty set of plausible game models under potential model mismatches. The resulting robust equilibrium provides a performance guarantee across all models within this set, thereby ensuring each agent’s resilience against model mismatch.
@@
The nature of a DR-MG problem is fundamentally shaped by the optimality criterion used. Much of the prior work has concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) or discounted-reward setting (zhang2020robust; kardecs2011discounted), where the objective is to maximize the expected reward over a finite number of steps or a discounted sum of rewards.
@@
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics (krnjaic2024scalable), communication networks (qu2020scalable), autonomous vehicle coordination (peake2020multi), financial markets (vadori2024towards), or even peer-to-peer energy trading (qiu2021multi). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate (hernandez2000zero). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium (filar2012competitive); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions (puterman1994markov; atia2021steady). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
@@
Moreover, none of the methods for DR-MGs under the discounted reward or finite horizon can be extended to the average-reward setting. The existence of (non-stationary) Nash Equilibrium under finite horizon is derived through backward induction (blanchet2023double), which does not apply to infinite-horizon settings. The existence of stationary Nash Equilibrium under the discounted reward criterion is derived in (kardecs2011discounted), which however heavily relies on the uniqueness of the solution to the discounted robust Bellman equation (iyengar2005robust) and hence cannot be applied to the average reward setting. Another natural attempt is to use the discounted reward as an intermediate step to study the robust Nash Equilibrium under the average reward by setting the discount factor approaching to $1$. Although this limit method is effective in the single-agent setting (wang2023robust), it cannot solve the DR-MGs, due to the complicated structure (also see our detailed discussion in Section 3). To date, the literature on average-reward DR-MGs remains sparse, leaving fundamental questions about the structure of robust equilibria and the design of convergent solution algorithms largely unanswered.
@@ 1.1 Contributions
In this paper, we develop a comprehensive study of DR-MGs under the average-reward formulation. Our contributions are summarized as follows.
Solvability of the Robust Bellman Equation under Irreducible Setting: As a foundational step, we first revisit the single-agent distributionally robust average-reward MDPs with the irreducible assumption. Despite existing studies largely utilizing the robust Bellman equation (wang2023robust; wang2023model; wang2024robust; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample), its solvability is not proved, but rather implicitly assumed. We establish the well-posedness of the robust average-reward Bellman equation, study the structure of its solutions, and connect them with the optimal robust policies. This analysis provides structural insights and tools that are essential for solving multi-agent Markov games when long-term performance is of interest.
Existence and Algorithmic Identification of Robust Nash Equilibrium: We further introduce a formulation of induced distributionally robust MDPs to reduce the Nash Equilibrium identification in comprehensive multi-agent games to optimizing single-agent distributionally robust MDPs. Under the irreducibility setting, we utilize the solvability results we obtained to show a correspondence between the optimal robust policy and solutions of the Bellman equation, and prove the convexity and hemi-continuity of the best-response set, deriving the existence of a stationary Nash Equilibrium. We further study the more general weak communicating setting, where no such correspondence can be obtained. We hence instead study a proxy mapping which is a subset of the best-response set, and derive the NE based on it. These results address the fundamental open question, demonstrating that stable solutions exist in these complex, robust, and long-term horizon competitive environments.
@@ 1.2 Related Works
Non-robust Markov Games. Markov Games (littman1994markov) provide the foundational mathematical framework for multi-agent sequential decision-making. The majority of early work focused on the discounted-reward setting, where the existence of a stationary Nash Equilibrium was established in (shapley1953stochastic). The average-reward criterion, while more suitable for systems with long operational horizons, presents greater analytical challenges. Unlike the discounted case, a stationary NE is not guaranteed to exist in a general average-reward MG. Its existence has only been proven under additional structural assumptions on the dynamics of the game (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; filar2012competitive).
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, distributionally robust MDPs are first developed in (iyengar2005robust; nilim2004robustness) for both finite and discounted settings. The studies under the average reward setting are recently developed in (wang2023robust; wang2023model; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample; grand2023beyond; grand2023reducing; chatterjee2023solving; wang2024robust). While the robust Bellman equation is central to these methods, its solvability was often implicitly assumed without rigorous proof.
Distributionally Robust Markov Games. Extending distributional robustness to the multi-agent setting is a recent and active area of research. However, the literature has overwhelmingly concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) and discounted-reward criteria (zhang2020robust; kardecs2011discounted). These settings are often more analytically tractable because the influence of future rewards diminishes, simplifying the analysis. However, none of these methods can be extended to the average-reward setting, as we shall discuss later.
@@ 2 Preliminaries and Problem Formulation
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (littman1994markov), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\otimes_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}_{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the actions of all agents. At some state $s$, all agents act simultaneously by taking joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
@@
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on stationary Nash Equilibrium (NE). More specifically, a stationary NE is a product policy $\pi^{\star}=(\pi^{\star}_{1},...,\pi^{\star}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{\star}$:
$V^{\pi^{\star}}_{\mathsf{P},i}(\rho)\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathsf{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}\in\Pi_{i},$ (1)
@@
for some initial distribution $\rho$ and $V^{\pi}_{\mathsf{P},i}(\rho)\triangleq\mathbb{E}_{s\sim\rho}[V^{\pi}_{\mathsf{P},i}(s)]$, and $\Pi_{i}$ denotes the set of stationary policies for agent $i$, $\pi^{\star}_{-i}$ denotes the joint policy of all other agents except $i$ from $\pi^{\star}$, and $(\pi^{\star}_{-i},\pi_{i})$ is the joint policy. The existence of stationary NE is derived for MGs with discounted reward (shapley1953stochastic) and average-reward MGs with additional structural assumptions (filar2012competitive).
@@
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\otimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ (iyengar2005robust; nilim2004robustness; zhang2020robust; blanchet2023double; shi2024sample), where the uncertainty set is independently defined over all $(s,a)$ pairs.
@@
After all agents take their actions $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}(\rho)$, and robust average reward $g^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}(\rho)$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{\star}$ such that for some initial state distribution $\rho$,
$g^{\pi^{\star}}_{\mathcal{P},i}(\rho)\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}.$ (2)